Grade 12Algebra

Find the product, using suitable properties:
15 × (–25) × (– 4) × (–10)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

The given equation is in the form of Commutative law of Multiplication.
=a × b=b ×a
Then,
= 15 × [(–25) × (–4)] × (–10)
= 15 × [100] × (–10)
= 15 × [-1000]
= - 15000

Thanks